Calculate the after-tax cost of the debt under each of the following conditions: rd of 9%,...

  1. Calculate the after-tax cost of the debt under each of the following conditions:
    1. rd of 9%, tax rate of 8%
    2. rd of 9%, tax rate of 28%
    3. rd of 9%, tax rate of 38%
Answer #1

After-tax cost of debt = Before-tax cost of debt * (1 - tax rate)

a. r = 0.09 * (1 - 0.08)

r = 0.0828

r = 8.28%

b. r = 0.09 * (1 - 0.28)

r = 0.0648

r = 6.48%

c. r = 0.09 * (1 - 0.38)

r = 0.0558

r = 5.58%

Can you please upvote? Thank You :-)